# A Survey on Future Frame Synthesis: Bridging Deterministic and Generative Approaches

### v7

@@ Abstract
Future Frame Synthesis (FFS) is dedicated to generating future frame sequences based on existing content, expanding the scope of Video Frame Prediction by emphasizing the synthesis aspect. This survey provides a comprehensive review of existing research on FFS, covering commonly used datasets and representative algorithms. We discuss key challenges and trace the evolution of FFS in computer vision, particularly the shift from deterministic to generative approaches. Our taxonomy outlines major advances and methodological shifts, emphasizing the rising significance of generative models in producing realistic and diverse predictions. This survey offers a comprehensive analysis of current research and, moreover, suggests promising avenues for future exploration in this ever-changing domain.
@@ 1 Introduction
The field of FFS is rapidly moving from traditional deterministic prediction to large-scale generative approaches as the latest video generative models continue to emerge. These models have made important advances in parameter scales, generation lengths, control capabilities, and training strategies, injecting new vitality. As a foundation, we introduce the problem formulation and core challenges in Section 2. Our taxonomy is organized around the degree of stochasticity in modeling approaches. In Table 2, we present deterministic algorithms that aim to perform pixel-level fitting based on fixed target frames. However, pixel-level metrics tend to drive models to average over multiple plausible futures, often resulting in blurry predictions. In Section 4, we examine algorithms that enable stochastic motion prediction. These include approaches that inject stochastic variables into deterministic models, as well as methods based on explicit probabilistic modeling. Such methods enable sampling from motion distributions, yielding diverse yet plausible predictions beyond the target frame. Given current FFS algorithms’ limited generative capacity, especially for high-resolution videos involving object appearance and disappearance, we discuss the generative FFS methods in Section 5. These methods prioritize producing coherent long-term sequences over pixel-level accuracy. In Section 6, we explore the broad applicability of FFS in areas such as world model, autonomous driving, robotics, film production, meteorology, and anomaly detection. These use cases demonstrate the role of FFS in dynamic scene understanding and interaction. In Section 7, we review prior surveys on video prediction and diffusion-based video generation. We also clarify our distinct focus: a comprehensive analysis of FFS spanning from deterministic to generative paradigms, emphasizing the growing role of generative models in producing realistic and diverse future predictions.
@@ 2.2 Paradigms and Architectures
In the domain of FFS, three paradigms have emerged—deterministic, stochastic, and generative—each representing a distinct modeling approach.
@@
The deterministic paradigm emphasizes pixel-level fitting to fixed target frames, typically employing low-level computer vision architectures such as Convolutional Neural Networks (CNNs) (Krizhevsky et al., 2012), Recurrent Neural Networks (RNNs) (Rumelhart et al., 1985), Long Short-Term Memory (LSTM) (Hochreiter & Schmidhuber, 1997) and U-Net (Ronneberger et al., 2015). Recently, the Transformer architecture has gradually challenged the dominance of traditional CNNs and RNNs. A landmark work, Vision Transformer (ViT) (Dosovitskiy et al., 2021), divides images into fixed-size patches and treats these patches as sequences input into the Transformer for processing. Swin Transformer (Liu et al., 2021) further introduces a hierarchical Transformer structure that handles features on different scales through a local window self-attention mechanism and progressively expanding receptive fields. In low-level vision tasks, IPT (Chen et al., 2021) employs pre-trained Transformers to address various low-level visual tasks. And Models like TimeSformer (Bertasius et al., 2021) and ViViT (Arnab et al., 2021) utilize Transformers to process spatio-temporal information in videos by combining video frames and time steps.
@@
For deterministic models, optimizing pixel-level metrics such as Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity (SSIM) (Wang et al., 2004) often leads to blurry outputs, as the models tend to average over multiple plausible futures.  Regarding the evaluation metrics, we will delve into the discussion in the next subsection. Unlike deterministic models that usually optimize pixel-level metrics, the stochastic paradigm introduces randomness into the prediction process by incorporating stochastic variables or distributions to model the inherent uncertainty in video dynamics. This approach aims to capture the variability and unpredictability of video sequences, often producing results that deviate significantly from the ground truth. Probabilistic models, such as Variational Autoencoders (VAEs) (Kingma & Welling, 2014) and Generative Adversarial Networks (GANs) (Goodfellow et al., 2020), are commonly used to achieve this. This approach does not necessarily aim to generate entirely new video content but rather to capture the variability in future outcomes. The generative paradigm, on the other hand, prioritizes the synthesis of coherent and plausible video sequences over pixel-level fidelity. It leverages advanced generative models, such as diffusion models and large language models (LLMs), to produce diverse and imaginative future frames that capture complex scene dynamics, including object emergence and disappearance. Recently, diffusion models have emerged as a powerful paradigm for generative tasks, including FFS. These models, such as DDPM (Ho et al., 2020), learn to generate data by progressively denoising a random noise signal. Diffusion models like Video Diffusion (Ho et al., 2022) and SVD (Blattmann et al., 2023a) have demonstrated remarkable ability to produce high-quality, diverse, and temporally coherent video sequences. By leveraging the gradual denoising process, these models can effectively capture complex scene dynamics and generate realistic future frames, even for long-term predictions. Furthermore, flow-matching models (Lipman et al., 2022; Dao et al., 2023) and rectified flow (Liu et al., 2022) have gained attention as an alternative approach to generative modeling, offering improved efficiency and scalability compared to traditional diffusion models. As research evolves, the boundaries between these paradigms continue to blur, with growing efforts to integrate their strengths and build more capable and versatile FFS systems.
@@ 3.1.1 Recurrent networks
PredNet (Lotter et al., 2017) pioneered the exploration of recurrent neural networks in video synthesis, drawing inspiration from predictive coding in neuroscience and employing a recurrent convolutional network to process video features effectively. Building on this foundation, PredRNN (Wang et al., 2017) introduces significant improvements by modifying LSTM architecture with a dual-memory structure, aiming to enhance spatio-temporal modeling. Despite these advances, the model still faces challenges such as gradient vanishing in video synthesis tasks. To address these limitations, ConvLSTM (Shi et al., 2015) emerges as a pivotal model by ingeniously integrating LSTM with CNNs to effectively capture motion and spatio-temporal dynamics—a development that has significantly influenced subsequent video synthesis models. E3d-LSTM (Wang et al., 2019) further advances the field by incorporating 3D convolutions into RNNs and introducing a gate-controlled self-attention module, thereby significantly improving long-term synthesis capabilities. However, the increased computational complexity introduced by 3D convolutions may offset the performance gains in certain applications. MSPred (Villar-Corrales et al., 2022) proposes a hierarchical convolutional-recurrent network that operates at multiple temporal frequencies to predict future video frames, as well as other representations such as poses and semantics.
@@ 3.1.3 Transformers
Since the groundbreaking design of ViT (Dosovitskiy et al., 2021), which applied a pure Transformer directly to sequences of image patches to extract continuous tokens, the use of Transformers in frame synthesis has attracted significant attention. Video frame interpolation is a task that is closely related to FFS (Liu et al., 2017). Shi et al. (2022) and Lu et al. (2022) propose Transformer-based video interpolation frameworks that leverage self-attention mechanisms to capture long-range dependencies and enhance content awareness. The primary difference from convolution is that the self-attention mechanism can dynamically compute the correlation between each patch and all other patches, thereby capturing global information, whereas convolution extracts features within local regions using fixed kernels. These methods further introduce innovative strategies, such as local attention in the spatio-temporal domain and cross-scale window-based attention, to improve performance and effectively handle large motions. Ye & Bilodeau (2023) present an efficient Transformer model for video prediction, leveraging a novel local spatio-temporal separation attention mechanism, and compares three variants: fully auto-regressive, partially auto-regressive, and non-auto-regressive, to balance performance and complexity. Numerous ongoing studies continue to explore the use of Transformers for modeling inter-frame dynamics under varying motion amplitudes and for addressing challenges in high-resolution frame synthesis (Park et al., 2023; Zhang et al., 2024).
@@
Many researchers believe that Transformers perform particularly well on large-scale datasets (Zhai et al., 2022; Smith et al., 2023). However, most of the existing studies focus primarily on high-level vision tasks. Furthermore, the success of Transformers in many tasks depends heavily on fully leveraging the capabilities of foundation models. In the context of image synthesis, best practices remain unclear (Li et al., 2023). Drawing insights from LLMs and related architectures may offer a promising direction, which we further explore in Section 5.2.
@@ 3.1.4 Motion-based synthesis
There are still ample opportunities for exploration in the representation of motion. For example, GaussianPrediction Zhao et al. (2024) explores the use of 3D Gaussian representations for modeling the appearance and geometry of dynamic scenes, enabling the rendering of images for future scenarios. MoSca (Lei et al., 2024) proposes the Motion Scaffold representation, which lifts the predictions of 2D foundational models (such as depth estimation, pixel trajectories, and optical flow) to 3D and optimizes them using physics-inspired constraints. Ultimately, MoSca reconstructs the geometry and appearance of the scene with a set of dynamic Gaussians (Wu et al., 2024b). Building on video diffusion models, V3D (Chen et al., 2024g) leverages their learned world simulation capabilities to perceive the 3D world, introduces a geometrical consistency prior, and finetunes the video diffusion model into a multi-view consistent 3D generator. Seamlessly integrating new representation methods with existing frameworks presents novel challenges.
@@ 8 Conclusion
In this survey, we have examined various aspects of FFS, including widely used datasets, evolving algorithmic paradigms, and prevailing challenges in the field.
Considering the broader development trends in artificial intelligence, we argue that the trajectory of video synthesis research should bifurcate. On the one hand, future research should focus on model lightweighting tailored for high-definition video applications, targeting low-level objectives such as video compression and short-term motion estimation. On the other hand, research should explore how models can develop a fundamental understanding of the physical world and generate content by leveraging substantial computational resources and diverse, long-duration video datasets. For the latter, future research should prioritize the development of evaluation metrics that incentivize stochastic synthesis, thereby expanding the potential for simulating the complexity of the human world. The ultimate objective is to develop video synthesis models with a profound understanding of the inherent dynamics within videos. Such models would be capable of generating videos over extended temporal horizons with high stochastic complexity in real-world scenarios.
Our proposed taxonomy is grounded in algorithmic stochasticity, highlighting the significant shift from deterministic approaches toward generative methodologies. This survey underscores the need to balance pixel-level accuracy with a deeper understanding of complex scene dynamics in video synthesis. Additionally, we examine the intricacies of stochastic emergence and disappearance processes, and advocate for improved evaluation metrics as well as the use of large-scale video datasets and substantial computational resources. We also categorize existing research directions and discuss the prevailing challenges in this domain. These insights are intended to inform and guide future research in video synthesis. As the field progresses, we anticipate the emergence of models with a more profound and nuanced understanding of real-world dynamics. Such advancements promise to improve accuracy, efficiency, and creative potential, paving the way for novel applications and future research.

### v8

@@ Abstract
Future Frame Synthesis (FFS), the task of generating subsequent video frames from context, represents a core challenge in machine intelligence and a cornerstone for developing predictive world models. This survey provides a comprehensive analysis of the FFS landscape, charting its critical evolution from deterministic algorithms focused on pixel-level accuracy to modern generative paradigms that prioritize semantic coherence and dynamic plausibility. We introduce a novel taxonomy organized by algorithmic stochasticity, which not only categorizes existing methods but also reveals the fundamental drivers—advances in architectures, datasets, and computational scale—behind this paradigm shift. Critically, our analysis identifies a bifurcation in the field’s trajectory: one path toward efficient, real-time prediction, and another toward large-scale, generative world simulation. By pinpointing key challenges and proposing concrete research questions for both frontiers, this survey serves as an essential guide for researchers aiming to advance the frontiers of visual dynamic modeling.
@@ 1 Introduction
Figure 1: Structure of our taxonomy. We categorize future frame synthesis (FFS) approaches based on the degree of stochasticity in the modeling paradigm.
The field of FFS is rapidly moving from traditional deterministic prediction to large-scale generative approaches as the latest video generative models continue to emerge. These models have made important advances in parameter scales, generation lengths, control capabilities, and training strategies, injecting new vitality. As a foundation, we introduce the problem formulation and core challenges in Section 2. Our taxonomy is organized around the degree of stochasticity in modeling approaches. In Table 2, we present deterministic algorithms that aim to perform pixel-level fitting based on fixed target frames. However, pixel-level metrics tend to drive models to average over multiple plausible futures, often resulting in blurry predictions. In Section 4, we examine algorithms that enable stochastic motion prediction. These include approaches that inject stochastic variables into deterministic models, as well as methods based on explicit probabilistic modeling. Such methods enable sampling from motion distributions, yielding diverse yet plausible predictions beyond the target frame. Given current FFS algorithms’ limited generative capacity, especially for high-resolution videos involving object appearance and disappearance, we discuss the generative FFS methods in Section 5. These methods prioritize producing coherent long-term sequences over pixel-level accuracy. Figure 1 visually illustrates the taxonomy and serves as a structural guide for the subsequent sections.
In Section 6, we explore the broad applicability of FFS in areas such as world model, autonomous driving, robotics, film production, meteorology, and anomaly detection. These use cases demonstrate the role of FFS in dynamic scene understanding and interaction. In Section 7, we review prior surveys on video prediction and diffusion-based video generation. We also clarify our distinct focus: a comprehensive analysis of FFS spanning from deterministic to generative paradigms, emphasizing the growing role of generative models in producing realistic and diverse future predictions.
@@ 2.2 Paradigms and architectures
In the domain of FFS, three paradigms have emerged—deterministic, stochastic, and generative—each representing a distinct modeling approach. Early deterministic methods are designed to minimize pixel-wise error. This objective inevitably produces blurry averages, which in turn motivate stochastic formulations that explicitly model uncertainty. Furthermore, the need for long-horizon and high-level plausibility pushes the field toward fully generative paradigms.
@@
The deterministic paradigm emphasizes pixel-level fitting to fixed target frames, typically employing low-level computer vision architectures such as Convolutional Neural Networks (CNNs) (Krizhevsky et al., 2012), Recurrent Neural Networks (RNNs) (Rumelhart et al., 1985), Long Short-Term Memory (LSTM) (Hochreiter & Schmidhuber, 1997) and U-Net (Ronneberger et al., 2015). Recently, the Transformer architecture has gradually challenged the dominance of traditional CNNs and RNNs. A landmark work, Vision Transformer (ViT) (Dosovitskiy et al., 2021), divides images into fixed-size patches and treats these patches as sequences input into the Transformer for processing. Swin Transformer (Liu et al., 2021) further introduces a hierarchical Transformer structure that handles features on different scales through a local window self-attention mechanism and progressively expanding receptive fields. In low-level vision tasks, IPT (Chen et al., 2021) employs pre-trained Transformers to address various low-level visual tasks. And Models like TimeSformer (Bertasius et al., 2021) and ViViT (Arnab et al., 2021) utilize Transformers to process spatio-temporal information in videos by combining video frames and time steps. During this period, models generally pursue efficiency and smaller parameter scales. This is constrained by the prevailing GPU computing power at the time and aligns with their primary goal of addressing short-term, low-resolution prediction tasks.
@@
For deterministic models, optimizing pixel-level metrics such as Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity (SSIM) (Wang et al., 2004) often leads to blurry outputs, as the models tend to average over multiple plausible futures.  Regarding the evaluation metrics, we will delve into the discussion in the next subsection. Unlike deterministic models that usually optimize pixel-level metrics, the stochastic paradigm introduces randomness into the prediction process by incorporating stochastic variables or distributions to model the inherent uncertainty in video dynamics. This approach aims to capture the variability and unpredictability of video sequences, often producing results that deviate significantly from the ground truth. Probabilistic models, such as Variational Autoencoders (VAEs) (Kingma & Welling, 2014) and Generative Adversarial Networks (GANs) (Goodfellow et al., 2020), are commonly used to achieve this. This approach does not necessarily aim to generate entirely new video content but rather to capture the variability in future outcomes. The generative paradigm, on the other hand, prioritizes the synthesis of coherent and plausible video sequences over pixel-level fidelity. It leverages advanced generative models, such as diffusion models and large language models (LLMs), to produce diverse and imaginative future frames that capture complex scene dynamics, including object emergence and disappearance. Recently, diffusion models have emerged as a powerful paradigm for generative tasks, including FFS. These models, such as DDPM (Ho et al., 2020), learn to generate data by progressively denoising a random noise signal. Diffusion models like Video Diffusion (Ho et al., 2022) and SVD (Blattmann et al., 2023a) have demonstrated remarkable ability to produce high-quality, diverse, and temporally coherent video sequences. By leveraging the gradual denoising process, these models can effectively capture complex scene dynamics and generate realistic future frames, even for long-term predictions. Furthermore, flow-matching models (Lipman et al., 2022; Dao et al., 2023) and rectified flow (Liu et al., 2022) have gained attention as an alternative approach to generative modeling, offering improved efficiency and scalability compared to traditional diffusion models. The rise of diffusion models and large auto-regressive models is inextricably linked to the significant growth in computational resources in recent years and the widespread utilization of large-scale, high-resolution web video datasets. As research evolves, the boundaries between these paradigms continue to blur, with growing efforts to integrate their strengths and build more capable and versatile FFS systems.
@@ 3.1.1 Recurrent networks
Due to the inherent temporal nature of video data, RNNs became a natural choice for early future frame synthesis tasks. The core idea of such methods is to use RNN’s memory units to model the temporal dependencies between frames. However, the main challenge lies in how to effectively extend RNN’s ability to process one-dimensional sequences to high-dimensional spatio-temporal data like videos. PredNet (Lotter et al., 2017) pioneered the exploration of recurrent neural networks in video synthesis, drawing inspiration from predictive coding in neuroscience and employing a recurrent convolutional network to process video features effectively. Building on this foundation, PredRNN (Wang et al., 2017) introduces significant improvements by modifying LSTM architecture with a dual-memory structure, aiming to enhance spatio-temporal modeling. Despite these advances, the model still faces challenges such as gradient vanishing in video synthesis tasks. To address these limitations, ConvLSTM (Shi et al., 2015) emerges as a pivotal model by ingeniously integrating LSTM with CNNs to effectively capture motion and spatio-temporal dynamics—a development that has significantly influenced subsequent video synthesis models. E3d-LSTM (Wang et al., 2019) further advances the field by incorporating 3D convolutions into RNNs and introducing a gate-controlled self-attention module, thereby significantly improving long-term synthesis capabilities. However, the increased computational complexity introduced by 3D convolutions may offset the performance gains in certain applications. MSPred (Villar-Corrales et al., 2022) proposes a hierarchical convolutional-recurrent network that operates at multiple temporal frequencies to predict future video frames, as well as other representations such as poses and semantics.
@@ 3.1.3 Transformers
Since the groundbreaking design of ViT (Dosovitskiy et al., 2021), which applied a pure Transformer directly to sequences of image patches to extract continuous tokens, the use of Transformers in frame synthesis has attracted significant attention. Video frame interpolation is a task closely related to FFS (Liu et al., 2017). Shi et al. (2022) and Lu et al. (2022) propose Transformer-based video interpolation frameworks that leverage self-attention mechanisms to capture long-range dependencies and enhance content awareness. The primary difference from convolution is that the self-attention mechanism can dynamically compute the correlation between each patch and all other patches, thereby capturing global information, whereas convolution extracts features within local regions using fixed kernels. These methods further introduce innovative strategies, such as local attention in the spatio-temporal domain and cross-scale window-based attention, to improve performance and effectively handle large motions. Ye & Bilodeau (2023) present an efficient Transformer model for video prediction, leveraging a novel local spatio-temporal separation attention mechanism, and compares three variants: fully auto-regressive, partially auto-regressive, and non-auto-regressive, to balance performance and complexity. Numerous ongoing studies continue to explore the use of Transformers for modeling inter-frame dynamics under varying motion amplitudes and for addressing challenges in high-resolution frame synthesis (Park et al., 2023; Zhang et al., 2024).
@@
Many researchers believe that Transformers perform particularly well on large-scale datasets (Zhai et al., 2022; Smith et al., 2023). However, most of the existing studies focus primarily on high-level vision tasks. Furthermore, the success of Transformers in many tasks depends heavily on fully leveraging the capabilities of foundation models. In the context of image synthesis, best practices remain unclear (Li et al., 2023). These challenges also explain why generative methods, especially token-based methods that directly draw on LLM architectures and training paradigms, which we will explore in Section 5.2, have become a more attractive direction for exploration.
@@ 3.1.4 Motion-based synthesis
There are still ample opportunities for exploration in the representation of motion. For example, GaussianPrediction Zhao et al. (2024) explores the use of 3D Gaussian representations to model the appearance and geometry of dynamic scenes, enabling image rendering for future scenarios. MoSca (Lei et al., 2024) proposes the Motion Scaffold representation, which lifts the predictions of 2D foundational models (such as depth estimation, pixel trajectories, and optical flow) to 3D and optimizes them using physics-inspired constraints. Ultimately, MoSca reconstructs the geometry and appearance of the scene with a set of dynamic Gaussians (Wu et al., 2024b). Building on video diffusion models, V3D (Chen et al., 2024g) leverages their learned world simulation capabilities to perceive the 3D world, introduces a geometrical consistency prior, and finetunes the video diffusion model into a multi-view consistent 3D generator. Seamlessly integrating new representation methods with existing frameworks presents novel challenges.
@@ 8 Conclusion
In this survey, we have systematically charted the evolution of Future Frame Synthesis (FFS), tracing its journey from deterministic prediction to the era of large-scale generative models. Our taxonomy, organized by the degree of algorithmic stochasticity, not only categorizes the field but also illuminates the fundamental methodological shifts that have defined its progress. Looking forward, we argue that the trajectory of FFS is not monolithic but is bifurcating into two distinct, yet potentially synergistic, research frontiers, each with its own set of grand challenges and open questions.
The First Frontier: Pragmatic Real-Time Synthesis. This direction prioritizes efficiency, precision, and applicability for high-definition, low-latency tasks. Rather than pursuing general world knowledge, its goal is to perfect low-level objectives like video compression, frame interpolation, and short-term forecasting. Key open questions in this domain include:
Hybrid Model Architectures: Can we design lightweight, deterministic modules—for instance, advanced optical flow or motion field estimators—that can be efficiently "plugged into" large generative models? Such hybrids could leverage the generative prior for context while using the deterministic module to enforce short-term, pixel-perfect motion accuracy at a low computational cost.
Knowledge Distillation for Motion: How can the rich, implicit motion understanding learned by massive models like SVD (Blattmann et al., 2023a) or Sora (Brooks et al., 2024) be effectively distilled into compact, recurrent-free networks? Developing novel distillation techniques tailored for temporal dynamics, rather than just static image features, is critical for real-world deployment on edge devices.
The Second Frontier: Generative World Simulation. This path is more ambitious, aiming to develop models with a fundamental understanding of the physical world, capable of generating diverse, coherent, and long-duration videos by leveraging vast computational and data resources. This research directly contributes to the long-term vision of creating general-purpose world models. The most pressing challenges here are not just about scaling, but about fundamental model capabilities:
Bridging the Paradigm Gap: How do we unify the strengths of disparate synthesis paradigms? Current methods excel at either precise motion warping (motion-based synthesis) or creative content generation (diffusion models), but rarely both. A pivotal breakthrough would be a unified framework that generates novel objects and events while adhering strictly to the predicted trajectories of existing elements.
Solving the Tokenizer Bottleneck: Is the persistent quality gap between token-based generative models and their diffusion-based counterparts a fundamental limitation of processing visual data as discrete sequences, or is it an engineering problem of tokenizer design and training infrastructure? Exploring hybrid auto-regressive/diffusion backbones or developing novel, lookup-free quantization schemes could unlock the scalability of LLM-like architectures for high-fidelity video generation.
From Semantic to Physical Control: Current controllability is largely limited to semantic instructions like text or sparse inputs like strokes. The next leap is to enable fine-grained, quantitative control over the underlying physics of a scene. Can we instruct a model to generate a video where "the ball has a mass of 5kg and a friction coefficient of 0.3," and see a physically plausible result? This requires moving beyond pattern recognition to building genuine, albeit simplified, simulation engines.
Progress in this second frontier is fundamentally gated by the evolution of our evaluation metrics. Metrics must evolve to reward not just perceptual quality, but also long-term temporal coherence, causal reasoning, and adherence to physical principles. The ultimate objective is to transform FFS from a task of visual pattern continuation into one of dynamic world simulation. As we have detailed, achieving this vision requires a concerted effort across model architecture, data curation, and evaluation philosophy. We anticipate that the insights and categorization provided in this survey will serve as a valuable roadmap for researchers navigating this exciting and rapidly changing domain.